# Synthetic History: Evaluating Visual Representations of the Past in Diffusion Models

### v2

@@ Abstract
As Text-to-Image (TTI) diffusion models become increasingly influential in content creation, growing attention is being directed toward their societal and cultural implications. While prior research has primarily examined demographic and cultural biases, the ability of these models to accurately represent historical contexts remains largely underexplored. To address this gap, we introduce a benchmark for evaluating how TTI models depict historical contexts. The benchmark combines HistVis, a dataset of 30,000 synthetic images generated by three state-of-the-art diffusion models from carefully designed prompts covering universal human activities across multiple historical periods, with a reproducible evaluation protocol. We evaluate generated imagery across three key aspects: (1) Implicit Stylistic Associations: examining default visual styles associated with specific eras; (2) Historical Consistency: identifying anachronisms such as modern artifacts in pre-modern contexts; and (3) Demographic Representation: comparing generated racial and gender distributions against historically plausible baselines. Our findings reveal systematic inaccuracies in historically themed generated imagery, as TTI models frequently stereotype past eras by incorporating unstated stylistic cues, introduce anachronisms, and fail to reflect plausible demographic patterns. By providing a reproducible benchmark for historical representation in generated imagery, this work provides an initial step toward building more historically accurate TTI models.
@@ 1 Introduction
Existing approaches to evaluating historical accuracy in foundation models mainly focus on identifiable landmarks (Weyand et al., 2020), prominent historical figures (Lee et al., 2023), or faithful reproduction of archival imagery (Akbulut et al., 2024). However, such approaches overlook a more fundamental question: how do TTI models construct visual interpretations of the past? The widely publicized case of Google’s Gemini (Team et al., 2024), which generated racially diverse Nazi soldiers, referred to as the “Black Nazi Problem” (Jacobi & Sag, 2024), highlighted the challenges of maintaining historical accuracy in AI-generated imagery and sparked debates on bias mitigation. Although AI bias is often discussed in terms of human representation, it also extends to the visual characteristics of generated images, such as the tendency of TTI models to associate certain historical periods with specific styles. Offert (2023), for instance, describes the difficulty of generating color images of 1930s fascist parades without outputs defaulting to visual features of early Kodachrome or colorized black-and-white photography. Besides demographic and stylistic biases, TTI models often fail to preserve temporal coherence, resulting in anachronistic depictions of objects, settings, or behaviors, such as smartphones in 18th-century scenes. Evaluating the historical reasoning of TTI models therefore requires a multifaceted approach that considers not only factual accuracy, but also implicit demographic assumptions, stylistic defaults, and consistent temporal representations.
To address these challenges, we introduce HistVis, a dataset of 30,000 synthetic images generated from 100 curated prompts covering 20 universal human activities (e.g., art, music) across five centuries (17th–21st) and five decades of the 20th century. The dataset is produced using three open-source diffusion models: Stable Diffusion XL (SDXL) (Podell et al., 2023), Stable Diffusion 3 (SD3) (Esser et al., 2024), and FLUX.1 Schnell111https://huggingface.co/black-forest-labs/FLUX.1-schnell . Building on this dataset, we construct a benchmark that pairs HistVis with a reproducible evaluation protocol to assess how TTI models depict historical contexts along three dimensions: (1) Implicit Stylistic Associations, examining how models default to specific visual styles across time periods; (2) Historical Consistency, identifying anachronisms such as modern artifacts in pre-modern contexts; and (3) Demographic Representation, assessing whether portrayals of race and gender align with historically plausible patterns. Although our experiments focus on HistVis, the evaluation framework is model-agnostic and directly applicable to any other TTI model.
@@ 2 Related Work
One significant dimension through which these tensions become evident is in demographic bias, particularly along gender and racial lines, which remains among the most extensively studied aspects of cultural encoding in TTI models. Studies consistently show that these biases manifest as stereotypical associations, such as overrepresenting women in caregiving roles and men in technical professions, thereby reinforcing existing social hierarchies (Seshadri et al., 2023; Hall et al., 2023). Methodologically, these patterns have been identified through comparisons with real-world benchmarks such as labor statistics (Luccioni et al., 2024; Bianchi et al., 2023; Luo et al., 2024), embedding-based bias metrics (Mandal et al., 2023), and classifier-based audits using models like BLIP-2 to evaluate responses to neutral occupational prompts (Cho et al., 2023). Recent work reveals that these biases not only affect the portrayal of gender but also associated objects and image layouts (Wu et al., 2024).
Beyond demographic biases, TTI models struggle with cultural representations, often reinforcing dominant narratives while failing to capture regional nuances. Minor prompt modifications, such as replacing a Latin character with a Greek or Arabic homoglyph, can dramatically shift outputs toward stereotypical depictions (Struppek et al., 2023). Neutral prompts also tend to produce Western-centric imagery, reflecting a lack of region-specific authenticity (Ventura et al., 2023). Benchmarks like CUBE (Senthilkumar et al., 2024) and UCOGC (Zhang et al., 2024a) show that even culturally specific prompts (e.g., India’s gopuram) often fail to yield accurate representations. Moreover, omitting country names can skew outputs toward U.S. or Indian-centric outputs, highlighting the need for globally aware and context-sensitive TTI systems (Basu et al., 2023).
@@
Although some recent work started to address history-related tasks, these efforts tend to focus on recognition and factual alignment. For example, CENTURY (Akbulut et al., 2024) investigates how accurately multimodal models describe sensitive historical photographs, while HEIM (Lee et al., 2023) and the Google Landmarks Dataset v2 (Weyand et al., 2020) evaluate recognition of historical figures and landmarks. However, these existing benchmarks focus on factual correctness and identification rather than probing how models contextualize representations of the past.
@@ 4 Implicit Stylistic Associations
The first aspect of our benchmark examines whether TTI models default to stylistic cues for historical periods, even when such features are not specified in the prompt. Leveraging the structure of the HistVis dataset where the activity remains fixed and only the time period varies, we introduce a methodology to quantify the strength and consistency of these associations across models and periods, enabling systematic analysis of a phenomenon previously described only qualitatively (Offert, 2023).
@@ 4.1 Experimental Setup
To quantify implicit stylistic associations, we develope a style predictor using a curated WikiArt corpus222https://www.wikiart.org/ of public-domain and permissively licensed images, organized into five broad, visually distinct style classes: drawing, engraving, illustration, painting, and photography. Although not exhaustive, these categories were selected because they capture distinct stylistic modes that emerged consistently in our initial qualitative analysis of historically themed generations, providing a structured foundation for identifying patterns of stylistic bias in generated imagery. The dataset consists of 1,280 training and 320 validation images per class (6,400 / 1,600 total). Since WikiArt does not distinguish between b&w and color photography, we apply a post-processing step using the colorfulness metric of Hasler & Suesstrunk (2003), re-labeling photographs with a score as monochrome.
We benchmark multiple encoders by training a linear probe on each frozen backbone, including CNNs (VGG-16, ResNet-50), vision transformers (Swin-Base, BEiT-Base, MAE ViT-Base, DINOv2 ViT-B/14), and a zero-shot CLIP baseline (ViT-B/32). DINOv2 (Oquab et al., 2023) achieves the strongest performance on the validation set (Accuracy = 0.89, F1 = 0.87), outperforming all others, and is therefore adopted for all downstream analyses. Using this predictor, we assign a stylistic label to every HistVis generation, enabling systematic measurement of model-specific tendencies in historical representation. Detailed metrics and confusion matrices for all backbones are reported in Appendix D.
@@ 4.2 Visual Style Dominance (VSD) Score
Table 1 shows VSD scores and dominant styles across models and time periods. In the 17th and 18th centuries, SDXL exhibits a strong preference for engravings, while SD3 and FLUX.1 favor paintings. In modern periods, FLUX.1 and SD3 converge on photography, whereas SDXL shifts toward illustrations. All three models display strong monochrome dominance in early 20th-century decades (1910s–1950s). Most VSD scores are statistically robust; however, for a few periods (e.g., FLUX.1 in the 1950s), overlapping CI preclude reliable identification of a single dominant style. While this may reflect classification uncertainty, it could also capture genuine cultural transitions, such as the shift from monochrome to color media during that period..
@@ 5.2 Anachronism Frequency & Severity Scores
Figure 4 plots the top 15 anachronistic elements per model, ranked independently by frequency and severity, aggregated across all centuries. Each data point corresponds to an element proposed by the LLM and confirmed via VQA, with fuzzy matching applied to unify lexical variations. For example, clothing appears in about 2-5% of all anachronistic images, making it the most frequent item overall, yet it shows relatively low severity, indicating a scattered but recurring misrepresentation. By contrast, items like audio devices and ironing equipment occur less often but with higher severity: when these elements are proposed by the LLM as plausible anachronisms for a given prompt, they are almost always identified in the corresponding generated images. This behavior suggests that models often rely more on the prompt’s conceptual cues (e.g., “listening to music”) than on the intended historical context, revealing a structural lapse in historical conditioning. Furthermore, SD3 appears most prone to anachronisms, with 20% of its 19th-century and 25% of its 1930s images flagged, followed by FLUX.1 and SDXL, revealing varying model tendencies. Examples of high-severity anachronisms are shown in Figure 3, with detailed breakdowns in Appendix J.
@@
To evaluate the alignment between our anachronism detection method and human judgment, we conducted an anonymous user study on a public crowdsourcing platform with no geographical restrictions. Since SD3 exhibited the highest anachronism rate, we sampled 200 images per decade and century from this model (1,800 in total). Each image was annotated by three participants, producing 5,400 responses, of which 2,040 high-quality annotations from 234 reliable participants were retained after filtering based on predefined control questions. Inter-annotator agreement, measured by Fleiss’ , indicated substantial agreement. Using the majority vote as the human baseline, our ensemble method reached 75% agreement with human judgments, substantially above chance, and outperformed individual VLMs (72% for GPT-4o, 68% for LLaMA-3.2, 63% for Qwen2.5). Appendix L provides additional details on the interface, instructions, and data collection.
@@ 6 Demographic Representation
The third dimension of our benchmark evaluates whether the TTI models portray gender and race in ways that align with historically plausible demographics rather than currentumptions or model-specific biases. However, establishing such plausibility is challenging, since historical records are often incomplete, vary across regions, and reflect shifting social roles. To approximate expectations at scale, we rely on LLM-based estimates that synthesize contextual knowledge for each activity and period. These estimates are not substitutes for historical records or expert annotation, but serve as reference points for identifying major divergences, cases where model outputs reflect stereotypical assumptions about the prompt activity itself rather than the specified temporal condition. While this approach provides a practical and scalable baseline, it remains a coarse approximation of historical accuracy and inherits the limitations and biases of the underlying language model.
@@ 6.1 Methodology
To evaluate demographic representation in historically conditioned TTI images, we implement a three-step pipeline that compares generated demographics to historically plausible expectations: (1) extracting attributes from generated images with an automated face classifier, (2) estimating expected distributions with an LLM that synthesizes contextual knowledge for each activity–time pair, and (3) quantifying under- and overrepresentation through metric-based comparison.
@@
To evaluate whether generated demographics align with historically plausible expectations, we use LLM-based estimates conditioned on each activity–period pair . For each case, the model is prompted with: “Given the [activity] and the [historical period], estimate the plausible demographic breakdown (in percentages) for gender and race.” This yields baseline distributions and .
We validated four candidate LLMs, GPT-4o, Claude 3.5 Sonnet, LLaMA-3.2–11B, Qwen2.5 VL–7B, against Our World in Data (OWID) statistics on gender and continent-level trends, focusing on categories with coverage since the 19th century: Education (Ritchie et al., 2023), Agriculture (Roser, 2023), and Work & Collaboration (Ortiz-Ospina & Roser, 2023). Race estimates were excluded during validation and replaced with continent-level distributions to avoid simplified regional–racial mappings, while our main analyses continue to use race labels from the FairFace classifier.
@@
We propose two metrics, underrepresentation and overrepresentation, to quantify how closely TTI-generated images align with historically plausible demographic expectations. Inspired by fairness metrics in ML that quantify deviations from reference distribution (Wang et al., 2022; Hall et al., 2023; Bianchi et al., 2023), our approach adapts these to the historical domain using LLM-derived estimates rather than population statistics. For each activity–period pair and demographic category , we compute both metrics. Formally, let denote the proportion of demographic group observed in the generated images, and the corresponding estimate provided by the LLM. We define:
@@ 6.2 Demographic Alignment Evaluation
When comparing GPT-4o-derived demographic baselines with face classifier predictions across TTI models, we find frequent divergences from historically plausible patterns, with variation depending on the model. As shown in Figure 6(a), FLUX.1 tends to overrepresent male figures, even in categories like "cooking & dining," where GPT-4o estimates a predominantly female presence until the 20th century. This male bias persists across centuries, including in "work & collaboration," which GPT-4o anticipates becoming gender-balanced by the 21st century. Conversely, “education & learning” skews female in the 17th–18th centuries, although GPT-4o indicates male majorities during that period. White individuals are generally overrepresented across models, though this trend diminishes by the 21st century, aligning closer to historical expectations. However, specific categories (e.g., "religion & spirituality," "agriculture") show pronounced spikes in non-White representations, as illustrated in Figure 6(b) for FLUX.1. Similar trends are visible across other models and time periods, suggesting that these distributions may reflect correlations in the training data rather than historically grounded depictions. Additional figures with representation scores across all models, time periods, and activity categories are provided in Appendix˜Q.
7 Discussion and Limitations
Our analysis reveals that TTI models rely on limited stylistic encodings rather than nuanced understandings of historical periods. Each era is strongly tied to a specific visual style, resulting in one-dimensional portrayals of history. Notably, photorealistic depictions of people appear only from the 20th century onward, with only rare exceptions in FLUX.1 and SD3, suggesting that models reinforce learned associations rather than flexibly adapting to historical contexts, perpetuating the notion that realism is a modern trait. In addition, frequent anachronisms suggest that historical periods are not cleanly separated in the latent spaces of these models, since modern artifacts often emerge in pre-modern settings, undermining the reliability of TTI systems in education and cultural heritage contexts. Finally, demographic representations underscore a lack of historical awareness: some outputs depict historically underrepresented groups at implausibly high rates, whereas others reproduce contemporary biases. This inconsistency highlights a deeper challenge in bias mitigation: correcting past exclusions risks distorting historical accuracy, yet faithfully reproducing historical inequalities may reinforce exclusionary narratives. Striking a balance between these goals remains therefore an open challenge.

### v3

@@ Abstract
As Text-to-Image (TTI) diffusion models become increasingly influential in content creation, growing attention is being directed toward their societal and cultural implications. While prior research has primarily examined demographic and cultural biases, the ability of these models to accurately represent historical contexts remains largely underexplored. To address this gap, we introduce a benchmark for evaluating how TTI models depict historical contexts. The benchmark combines HistVis, a dataset of 30,000 synthetic images generated by three state-of-the-art diffusion models from carefully designed prompts covering universal human activities across multiple historical periods, with a reproducible evaluation protocol. We evaluate generated imagery across three key aspects: (1) Implicit Stylistic Associations: examining default visual styles associated with specific eras; (2) Historical Consistency: identifying anachronisms such as modern artifacts in pre-modern contexts; and (3) Demographic Representation: comparing generated racial and gender distributions against LLM-estimated historically plausible demographics. Our findings reveal systematic inaccuracies in historically themed generated imagery, as TTI models frequently stereotype past eras by incorporating unstated stylistic cues, introduce anachronisms, and fail to reflect plausible demographic patterns. By providing a reproducible benchmark for historical representation in generated imagery, this work provides an initial step toward building more historically accurate TTI models.
@@ 1 Introduction
Existing approaches to evaluating historical accuracy in foundation models mainly focus on identifiable landmarks (Weyand et al., 2020), prominent historical figures (Lee et al., 2023), or objective description of archival imagery (Akbulut et al., 2025). However, such approaches overlook a more fundamental question: how do TTI models construct visual interpretations of the past? The widely publicized case of Google’s Gemini (Team et al., 2024), which generated racially diverse Nazi soldiers, referred to as the “Black Nazi Problem” (Jacobi & Sag, 2024), highlighted the challenges of maintaining historical accuracy in AI-generated imagery and sparked debates on bias mitigation. Although AI bias is often discussed in terms of human representation, it also extends to the visual characteristics of generated images, such as the tendency of TTI models to associate certain historical periods with specific styles. Offert (2023), for instance, describes the difficulty of generating color images of 1930s fascist parades without outputs defaulting to visual features of early Kodachrome or colorized black-and-white photography. Besides demographic and stylistic biases, TTI models often fail to preserve chronological coherence, resulting in anachronistic depictions of objects, settings, or behaviors, such as smartphones in 18th-century scenes. Evaluating the historical reasoning of TTI models therefore requires a multifaceted approach that considers not only factual accuracy, but also implicit demographic assumptions, stylistic defaults, and consistent chronological representations.
To address these challenges, we introduce HistVis, a dataset of 30,000 synthetic images generated from 100 curated prompts covering 20 universal human activities (e.g., art, music) across five centuries (17th–21st) and five decades of the 20th century. The dataset is produced using three open-source diffusion models: Stable Diffusion XL (SDXL) (Podell et al., 2023), Stable Diffusion 3 (SD3) (Esser et al., 2024), and FLUX.1 Schnell (Black Forest Labs, 2025). Building on this dataset, we construct a benchmark that pairs HistVis with a reproducible evaluation protocol to assess how TTI models depict historical contexts along three dimensions: (1) Implicit Stylistic Associations, examining how models default to specific visual styles across time periods; (2) Historical Consistency, identifying anachronisms such as modern artifacts in pre-modern contexts; and (3) Demographic Representation, assessing whether portrayals of race and gender align with historically plausible patterns. Although our experiments focus on HistVis, the evaluation framework is model-agnostic and directly applicable to any other TTI model.
@@ 2 Related Work
One significant dimension through which these tensions become evident is in demographic bias, particularly along gender and racial lines, which remains among the most extensively studied aspects of cultural encoding in TTI models. Studies consistently show that these biases manifest as stereotypical associations, such as overrepresenting women in caregiving roles and men in technical professions, thereby reinforcing existing social hierarchies (Seshadri et al., 2023; Hall et al., 2023b). Methodologically, these patterns have been identified through comparisons with real-world benchmarks such as labor statistics (Luccioni et al., 2024; Bianchi et al., 2023; Luo et al., 2024), embedding-based bias metrics (Mandal et al., 2023), and classifier-based audits using models like BLIP-2 to evaluate responses to neutral occupational prompts (Cho et al., 2023). Recent work reveals that these biases not only affect the portrayal of gender but also associated objects and image layouts (Wu et al., 2024).
Beyond demographic biases, TTI models struggle with cultural representations, often reinforcing dominant narratives and missing regional nuance. Benchmarks such as CUBE Senthilkumar et al. (2024), UCOGC Zhang et al. (2024a), CulturalFrames (Nayak et al., 2025), and cross-cultural annotation studies (Hall et al., 2024) show that even culturally specific prompts (e.g., India’s gopuram) frequently yield inaccurate depictions. Minor prompt changes, such as replacing a Latin character with a Greek or Arabic homoglyph, can shift outputs toward stereotypes (Struppek et al., 2023), while neutral prompts often produce Western-centric imagery with limited regional authenticity (Ventura et al., 2023; Hall et al., 2023a). Also, omitting country names can skew outputs toward U.S.-centric imagery, highlighting the need for globally aware, context-sensitive TTI models (Basu et al., 2023).
@@
Although some recent work started to address history-related tasks, these efforts tend to focus on recognition and factual alignment. For example, CENTURY (Akbulut et al., 2025) investigates how accurately and objectively multimodal models describe sensitive historical photographs, while HEIM (Lee et al., 2023) and the Google Landmarks Dataset v2 (Weyand et al., 2020) evaluate recognition of historical figures and landmarks. However, these existing benchmarks focus on factual correctness and identification rather than probing how models contextualize representations of the past.
@@ 4 Implicit Stylistic Associations
The first aspect of our benchmark examines whether TTI models default to stylistic cues for historical periods, even when such features are not specified in the prompt. Although these implicit stylistic associations do not introduce factual errors, they can reinforce stereotypes and fixed assumptions about how different historical periods are “supposed” to look. Leveraging, therefore, the structure of the HistVis dataset where the activity remains fixed and only the time period varies, we introduce a methodology to quantify the strength and consistency of these associations across models and periods, enabling systematic analysis of a phenomenon previously described only qualitatively (Offert, 2023).
@@ 4.1 Experimental Setup
To quantify implicit stylistic associations, we develop a style predictor using curated WikiArt corpus (WikiArt, 2025) of public-domain and permissively licensed images, organized into five broad, visually distinct style classes: drawing, engraving, illustration, painting, and photography. Although not exhaustive, these categories were selected because they capture distinct stylistic modes that emerged consistently in our initial qualitative analysis of historically themed generations, providing a structured foundation for identifying patterns of stylistic bias in generated imagery. The dataset consists of 1,280 training and 320 validation images per class (6,400 / 1,600 total). Since WikiArt does not distinguish between b&w and color photography, we apply a post-processing step using the colorfulness metric of Hasler & Suesstrunk (2003), re-labeling photographs with a score as monochrome.
We benchmark multiple encoders by training a linear probe on each frozen backbone, including CNNs (VGG-16, ResNet-50), vision transformers (Swin-Base, BEiT-Base, MAE ViT-Base, DINOv2 ViT-B/14), and a zero-shot CLIP baseline (ViT-B/32). DINOv2 (Oquab et al., 2023) achieves the strongest performance on the validation set (Acc. = 0.89, F1 = 0.87), outperforming all others, and is therefore adopted for all downstream analyses. Using this predictor, we assign a stylistic label to every HistVis generation, enabling systematic measurement of model-specific tendencies in historical representation. Detailed metrics and confusion matrices for all backbones are reported in Appendix D.
@@ 4.2 Visual Style Dominance (VSD) Score
Table 1 shows VSD scores and dominant styles across models and time periods. In the 17th and 18th centuries, SDXL exhibits a strong preference for engravings, while SD3 and FLUX.1 favor paintings. In modern periods, FLUX.1 and SD3 converge on photography, whereas SDXL shifts toward illustrations. All three models display strong monochrome dominance in early 20th-century decades (1910s–1950s). Most VSD scores are statistically robust; however, for a few periods (e.g., FLUX.1 in the 1950s), overlapping CI preclude reliable identification of a single dominant style. While this may reflect classification uncertainty, it could also capture genuine cultural transitions, such as the shift from monochrome to color media during that period.
@@ 5.2 Anachronism Frequency & Severity Scores
Figure 4 plots the top 15 anachronistic elements per model, ranked independently by frequency and severity, aggregated across all centuries. Each data point corresponds to an element proposed by the LLM and confirmed via VQA, with fuzzy matching applied to unify lexical variations. For example, clothing appears in about 2-5% of all anachronistic images, making it the most frequent item overall, yet it shows relatively low severity, indicating a scattered but recurring misrepresentation. By contrast, items like audio devices and ironing equipment occur less often but with higher severity: when these elements are proposed by the LLM as plausible anachronisms for a given prompt, they are almost always identified in the corresponding generated images. This behavior suggests that models often rely more on the prompt’s conceptual cues (e.g., “listening to music”) than on the intended historical context, revealing a structural lapse in historical conditioning. Furthermore, S3 appears most prone to anachronisms, with 20% of its 19th-century and 25% of its 1930s images flagged, followed by FLUX.1 and SDXL, revealing varying model tendencies. Examples of high-severity anachronisms are shown in Figure 3, with detailed breakdowns in Appendix J.
@@
To evaluate the alignment between our anachronism detection method and human judgment, we conducted an anonymous user study on a public crowdsourcing platform with no geographical restrictions. Since SD3 exhibited the highest anachronism rate, we sampled 200 images per decade and century from this model (1,800 in total). Each image was annotated by three participants, producing 5,400 responses, of which 2,040 high-quality annotations from 234 reliable participants were retained after filtering based on predefined control questions. Inter-annotator agreement, measured by Fleiss’ , indicated substantial agreement. Using the majority vote as the human baseline, our ensemble method reached 75% agreement with human judgments, substantially above chance, and outperformed individual VLMs (72% for GPT-4o, 68% for LLaMA-3.2, 63% for Qwen2.5). For context, prior work (D’Incà et al., 2024) reports 67% agreement on simpler contemporary VQA tasks (e.g., identifying the color of a car), suggesting that our approach performs competitively in a substantially richer historical and semantic setting. Appendix L provides additional details on the interface, instructions, and data collection.
@@ 6 Demographic Representation
The third dimension of our benchmark evaluates how TTI models portray gender and race in relation to LLM-predicted historically plausible demographics. The goal of this analysis is to reveal whether TTI models exhibit extreme patterns of over- and under-representation of certain demographic groups with respect to specific activities and time periods. Because our benchmark and dataset target broad period categories and general activities, a factual historical baseline grounded in real-world data across the full range of cultures, regions, periods and activities cannot be reliably established within the scope of this work. However, to still address historical demographic bias at least at a coarse level, we extract LLM-based estimates that synthesize contextual knowledge for each activity and period as starting points for analyzing over- and under-representation. These estimates are not substitutes for historical records or expert annotation, but serve as reference points for identifying major divergences, cases where model outputs reflect stereotypical assumptions about the prompt activity itself rather than the specified temporal condition. While this approach provides a practical and scalable baseline, it remains a coarse approximation of historical accuracy and inherits the limitations and biases of the underlying language model.
@@ 6.1 Methodology
To examine demographic representation in historically conditioned TTI images, we implement a three-step pipeline that compares generated demographics to LLM-predicted historically plausible expectations: (1) extracting attributes from generated images with an automated face classifier, (2) estimating expected distributions with an LLM that synthesizes contextual knowledge for each activity–time pair, and (3) quantifying under- and overrepresentation through metric-based comparison.
@@
To generate estimates of historically plausible expectations, we use LLM-based predictions conditioned on each activity–period pair . For each case, the model is prompted with: “Given the [activity] and the [historical period], estimate the plausible demographic breakdown (in percentages) for gender and race.” This yields baseline distributions and . Although real-world data is not available for the full scope of our categories and periods, we partially assess the LLM-based estimates by comparing them with the closest available data for three categories. We validated four candidate LLMs, GPT-4o, Claude 3.5 Sonnet, LLaMA-3.2–11B, Qwen2.5 VL–7B, against Our World in Data (OWID) statistics on gender and continent-level trends, focusing on categories with coverage since the 19th century: Education (Ritchie et al., 2023), Agriculture (Roser, 2023), and Work & Collaboration (Ortiz-Ospina & Roser, 2023). Because OWID does not provide race-based statistics and to avoid oversimplified region–race mappings, this evaluation of LLMs replaces race estimation with continent-level estimation. Our main analyses compare LLM race estimates with predictions derived from the FairFace classifier.
@@
We propose two metrics, underrepresentation and overrepresentation, to quantify how closely TTI-generated images align with the estimated historically plausible demographic expectations. Inspired by fairness metrics in ML that quantify deviations from reference distribution (Wang et al., 2022; Hall et al., 2023b; Bianchi et al., 2023), our approach adapts these to the historical domain using LLM-derived estimates rather than population statistics. For each activity–period pair and demographic category , we compute both metrics. Formally, let denote the proportion of demographic group observed in the generated images, and the corresponding estimate provided by the LLM. We define:
@@ 6.2 Demographic Alignment Evaluation
When comparing GPT-4o-derived demographic baselines with face classifier predictions across TTI models, we find frequent divergences from the estimated historically plausible patterns, with variation depending on the model. As shown in Figure 6(a), FLUX.1 tends to overrepresent male figures, even in categories like "cooking & dining," where GPT-4o estimates a predominantly female presence until the 20th century. This male bias persists across centuries, including in "work & collaboration," which GPT-4o anticipates becoming gender-balanced by the 21st century. Conversely, “education & learning” skews female in the 17th–18th centuries, although GPT-4o indicates male majorities for that period. White individuals are generally overrepresented across models, though this trend diminishes by the 21st century, aligning closer to LLM-estimated historical expectations. However, specific categories (e.g., "religion & spirituality," "agriculture") show pronounced spikes in non-White representations, as illustrated in Figure 6(b) for FLUX.1. Similar trends are visible across other models and time periods, suggesting that these distributions may reflect correlations in the training data rather than historically grounded depictions. Additional figures with representation scores across all models, time periods, and activity categories are provided in Appendix˜Q.
7 Discussion
7.1 Limitations